OpenStudy (anonymous):

How to find the vertical asymptotes of the rational function g(x)=2x/x-8?

OpenStudy (anonymous):

Do yo know how to use limits?

OpenStudy (anonymous):

No I've never solved this type of problem.

OpenStudy (anonymous):

Well, i'll supose you do, if you have a function f(x), there's a vertical asymptote x=x1 if \[\lim_{x \rightarrow x1} f(x)= \infty\] So, in your case, you have to \[\lim_{x \rightarrow x1} \dfrac {2x} {x-8} = \infty\] Limits go to infinite when, for example their denominator goes to 0 Get it?

OpenStudy (anonymous):

So the answer choices are x=8; x=-8; x=2; none

OpenStudy (anonymous):

No not really.....

OpenStudy (anonymous):

Have you ever seen real function limits?

OpenStudy (anonymous):

No this is a problem our teacher gave us, but we've never gone over it.

OpenStudy (anonymous):

Aww, that turns it a little bit hard to explain, if you don't mind reading... http://www.mash.dept.shef.ac.uk/Resources/limits.pdf This explains what limits are- (It's precalculus stuff)

OpenStudy (anonymous):

Ok thanks.